OpenStudy (suwhitney):

Solve.. 8^x=1/2^(-3/x) Quick question...

OpenStudy (suwhitney):

\[8^x=\frac{ 1 }{ 2^{\frac{ -3 }{ x}} }\]

OpenStudy (suwhitney):

Ok, so when I am flipped the fraction does it become.. \[8^x=(2^3)^\frac{ 3 }{ x }\] ?

OpenStudy (anonymous):

\[8^{x}=2^{\frac{ 3 }{ x }}\]

OpenStudy (anonymous):

take the log of both sides to bring the exponent down. If you remember \[\log 10^{x}=x \log10\]

OpenStudy (anonymous):

\[x \log 8 = \frac{ 3 }{ x } \log 2\]\[x ^{2} = \frac{ 3 \log 2 }{ \log 8 }\]\[x = \pm \sqrt{\frac{ 3 \log 2 }{ \log 8 }}\]